# QUITE: A Query Rewrite System Beyond Rules with LLM Agents

### v2

@@ 1. Introduction
Inefficient SQL queries remain a long-standing challenge in many database applications, often arising from poorly written queries by inexperienced users or automatically generated systems (Li, 2019; Zhou et al., 2020; Pirahesh et al., 1992). Query rewrite is a promising approach that transforms an original SQL into a semantically equivalent form with better performance (Finance and Gardarin, 1991a; Pirahesh et al., 1992; Wang et al., 2022). An effective rewritten query can improve the execution time of a poorly written SQL query by several orders of magnitude (Wang et al., 2022; Zhou et al., 2021).
Existing query rewrite approaches typically rely on predefined rewrite rules (Begoli et al., 2018; Graefe and McKenna, 1993; PostgreSQL Global Development Group, 2024; Zhou et al., 2021, 2023; Bai et al., 2023; Li et al., 2024; Sun et al., 2024). Heuristic-based methods apply these rules in a fixed order derived from practical experience (Begoli et al., 2018; Graefe and McKenna, 1993; PostgreSQL Global Development Group, 2024). However, such fixed orders often fail to generalize across different environments and query patterns. Thus, machine learning (ML)-based methods train models on rewrite histories to predict which rules to use and the order in which to apply them (Zhou et al., 2021, 2023). Recently, some approaches have explored the use of Large Language Models (LLMs) to select rewrite rules and determine their application order based on domain knowledge (Li et al., 2024; Sun et al., 2024).
@@
However, the above approaches can handle only a small subset of queries and may lead to performance regressions (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024). This limitation stems from the intrinsic challenges of rule-based query rewrite. First, the rich features of SQL and the nuances of semantics make it rather difficult to discover and verify new rewrite rules (Pirahesh et al., 1992; Chu et al., 2017b; Wang et al., 2022). Even state-of-the-art rule discovery system (Wang et al., 2022) can only support a narrow subset of operators. Second, fixed rewrite rules rely on pattern matching and therefore are fundamentally unable to optimize unseen or complex query patterns (Leis et al., 2015; Bai et al., 2023; Dong et al., 2023). Third, many effective rewrite strategies cannot be expressed by rewrite rules (Dong et al., 2023). For example, the Common Table Expressions (CTE) conversion in Example 1.1 cannot be captured by rewrite rules. This is because such conversion requires explicit control over evaluation order and sub-plan sharing, which is beyond the scope of relational-algebra rewrite rules (Graefe, 1993; Wang et al., 2022; Begoli et al., 2018).
Given these limitations, we revisit the query rewrite problem from the perspective of human experts and identify three key capabilities that humans exhibit, but rule-based systems lack. (1) Humans can perform rewrites that follow rules but go beyond the expressiveness of pattern matching (Example 1.1). (2) Humans can conduct context-aware rewrites by considering the underlying data distributions and cost estimates provided by the query optimizer (Example 1.2). (3) Experts can apply intent-oriented rewrites that capture the underlying goal of the query, which are not easily expressed as explicit rules (Example 1.3).
@@
Query Rewrite Beyond Rewrite Rules. Figure 1 (a) is a simplified query from the DSB benchmark (Ding et al., 2021). To avoid repeated calculation of nested subqueries, Figure 1 (b) rewrites them using Common Table Expressions (CTEs). This enables the database engine to compute intermediate results once and reuse them, leading to a 37.5 $\times$ reduction in query latency in our experimental setup.
@@
Context-Aware Query Rewrite. In Figure 1 (c), instead of directly joining two large tables (the sale with 10 M rows and the ttime with 1.4 M rows), we first joins the sale table with the prod table (100 K rows). This highly selective join would produce an intermediate result of only  100 K rows. Next we join this with ttime table and yield a 2.79 $\times$ speedup over (b) in our experimental setup. Such rewrites, which reduce intermediate size, are rare in rule-based systems and require analyzing data distribution and selectivity.
@@
Query Rewrite based on Query Intent. Figure 2 (a) shows a real-world query from Apache Calcite (apa, 2021). Rule-based rewrite cannot optimize this query (Figure 2 (b)). Identifying that the query intent is to sum salaries directly from the EMP table, human can simplify the original query by eliminating unnecessary joins and sub-queries that do not contribute to the intent of the query, achieving a 2.17 $\times$ execution improvement in our experimental setup.
@@
Although human experts can produce significantly better SQL rewrites than rule-based methods, manual rewrite suffers from poor scalability, especially when there are thousands or even more poorly written queries to optimize in cloud environments. Recently, LLMs have demonstrated nearly human-level semantic and reasoning abilities, making them a natural choice for query rewrite. However, this task remains challenging due to the hallucination problem of LLMs: they may generate responses that are plausible but factually incorrect or semantically invalid. Specifically, there are two fundamental challenges. C1. Ensuring Equivalent Rewrites. Producing equivalent rewrites requires a precise understanding of the rich features of SQL and the nuances of semantics. Even the most powerful LLMs produce rewrites with syntax or semantic errors in nearly 20% of cases (Section 7). C2. Ensuring Optimized Rewrites. Determining whether a rewritten query improves performance depends on multiple factors, including query intent, data distribution, optimizer estimates, system states, index configurations, and others. These contextual factors are typically unavailable to LLMs, and enabling LLMs to effectively utilize such information remains a non-trivial problem (Lao et al., 2024a; Bai et al., 2024; Tan et al., 2025).
@@
To address the limitations of rule-based query rewrite, the poor scalability of manual rewrite, and the unreliable nature of LLMs, we propose QUITE, a training-free and feedback-aware system that smartly leverages LLMs to rewrite SQL queries into semantically equivalent forms with significantly improved performance. QUITE supports a broader range of query patterns and rewrite strategies, while maintaining good scalability. First, to enable LLMs to capture query data characteristics and leverage database feedback, we design a multi-agent framework controlled by a finite state machine (FSM) to iteratively refine the rewrite process. We decompose the complex rewrite process into a set of subtasks, each handled by a specialized LLM agent to reduce hallucinations and improve rewrite ability (addressing C1, C2). Second, to enhance the ability of LLMs to generate high-quality query rewrites, we provide a rewrite middleware that includes: (1) a structured knowledge base extracted from official documentation and query rewrite discussions on web forums, (2) a hybrid SQL corrector to ensure query equivalence, and (3) an agent memory buffer that manages important rewrite contexts (addressing C1, C2). Third, we employ a novel hint injection technique to generate better execution plans for rewritten queries. This ensures that query execution exactly follows the optimized plan recommended by LLM agents, avoiding unintended modifications by the query optimizer due to inaccurate cost estimations. Additionally, this technique enables further improvements by incorporating physical-level optimizations (addressing C2).
@@
We extensively evaluate QUITE against state-of-the-art query rewrite systems on widely used benchmarks (e.g., TPC-H, DSB, and Calcite) and show that our system delivers up to a 21.9% speedup in query execution performance over the best-performing alternative. Moreover, it delivers 24.1% additional rewrites compared to that alternative, covering query cases that previous systems could not handle. In summary, we make the following contributions:
@@
  • We design a multi-agent framework controlled by a finite state machine to enable LLMs to use external tools and enhance the rewrite process with database feedback (Section 4).
@@ 2.1. Query Rewrite
(1) Discovery of new rules. Approaches such as WeTune (Wang et al., 2022) automatically generate and verify logical plan transformations, but they are limited to specific operator types and algebraic equivalences. Equality saturation (Nandi et al., 2021) and user-guided tools like QueryBooster (Bai et al., 2023) lack full autonomy and broad applicability. Recently, GenRewrite (Liu and Mozafari, 2024) uses LLMs to suggest rules in natural language. However, it still struggles with optimal rewrite selection because executing all candidates to measure latency incurs prohibitive overhead and its off-the-shelf verifiers permit equivalence checking for only a subset of rewrite pairs (Liu and Mozafari, 2024; Dong et al., 2023).
(2) Effective use of existing rules. A rule-based rewrite framework must answer two key questions: which rules to include in the rule set and how to select among applicable rules during rewriting. LearnedRewrite (Zhou et al., 2021) uses Monte Carlo Tree Search (MCTS) with learned cost models to explore the space of rule applications, but its performance depends on the accuracy of cost model estimation. LLM-R2 (Li et al., 2024) uses a trained selection model to retrieve the past rewrite query and rule demonstrations similar to the input SQL, and leverages them as an in-context prompt for the LLM to select promising rewrite rules. However, it still incurs high costs by executing every query demonstration on each rule set and faces distribution shifts between the demonstration and input SQL, since structurally similar queries may differ in row count, leading to suboptimal rule selection.
@@ 2.2. Query Hints
Query hints are SQL extensions that provide instructions to the database’s query engine to influence the selection of execution plans (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021). These hints allow users to exert precise control over how the database optimizer chooses physical operators and access methods.
As shown below for pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012), the hint /*+ HashJoin(employees departments) */ forces the query engine to use a hash join for the $employees\bowtie departments$ operation, while the /*+ SeqScan(employees)*/ enforces a full table scan of the employees relation. These hints can override the optimizer’s default cost-based plan selection.
HashJoin(employees departments)
SeqScan(employees)
@@
Although existing studies on query hints (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023) have shown promising results, they typically focus on boolean hints applied at the query level (e.g., the enable_nestloop configuration parameter in PostgreSQL), which apply globally and can affect subsequent workload queries. In contrast, pg_hint_plan injects fine‐grained hints at the level of individual joins or cardinality estimates, allowing the use of multiple hint types (e.g., cardinality adjustments) at a more fine-grained level (e.g., specifying join algorithms for particular table pairs) while requiring a more precise selection mechanism.
@@ 2.3. LLM and LLM Agents
LLM for Database Tasks. LLMs are increasingly applied to database tasks like text-to-SQL (Gao et al., 2023), SQL workload generation (Lao and Trummer, 2025), data processing (Albalak et al., 2024), and database optimization (Lao et al., 2024a; Zhao et al., 2025; Lao et al., 2024b; Tan et al., 2025; Lao et al., 2025), with recent work focusing on query optimization (Li et al., 2024; Sun et al., 2024; Liu and Mozafari, 2024). Although these methods show that LLMs can understand and modify SQL, the inherent complexity of query rewrite presents substantial challenges.
LLM Agents. The primary advantage of LLM agents over LLMs is their ability to collect feedback from the execution environment (e.g., database execution metrics) and dadapt their subsequent reasoning and outputs accordingly (Wu et al., 2025). They have proven particularly effective in complex problem solving—for example, SWE‐agent in software engineering (Yang et al., 2024) and ChemCrow in chemistry (Bran et al., 2023). QUITE applies this LLM agents methodology to query rewrite by gathering DBMS feedback and adjusting each query candidate accordingly. This performs a feedback-driven refinement, enabling reliable and advanced query rewrite that extends beyond predefined rules or the capabilities of standalone LLMs.
2.4. Comparison with Other Works
Instead of focusing on finding optimal rewrite rules with limited coverage (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024), QUITE addresses the core aspects of query rewriting by identifying semantic and physical rewrite bottlenecks through real-time interaction between LLM agents and the database, enhancing the detection of performance issues in the input SQL. Furthermore, QUITE generates a reasoning chain with natural language rewrite proposals and rewrites the SQL through multi-agent collaboration within a carefully designed Finite State Machine (FSM) pipeline and agent tools, ensuring equivalence and effectiveness. This approach enhances adaptability and efficiency in context-aware rewrites, compared to rule-based systems that miss complex query patterns and context-specific optimizations.
In contrast to GenRewrite (Liu and Mozafari, 2024), which directly utilizes LLMs to generate rewrite rules for rewriting, QUITE considers key factors such as the SQL’s physical plan, the database distribution, and the domain expert knowledge for analysis, ensuring better alignment with the execution environment. Additionally, QUITE introduces a specialized reasoning agent, which forms the core of its multi-agent framework, incorporating a planning mechanism to sequence high-impact rewrite actions, maximizing performance improvements and preventing inefficiencies and regressions.
Furthermore, QUITE incorporates a fine-grained hint injection technique into query rewrite process to further improve the query execution performance. To our knowledge, no other query rewrite systems have supported for these components.
@@ 3. System Overview
We present the system architecture of QUITE  as shown in Figure 3. QUITE comprises three components: LLM Agent-Based Query Rewriter, Rewrite Middleware, and Query Hint Recommender.
LLM Agent-Based Query Rewriter. The rewriter uses a training-free and feedback-aware LLM agent-based FSM to rewrite queries in multiple stages. First, ❶ a user provides the input SQL along with the corresponding database configuration (e.g., DBMS type, host address). Second, ❷ The rewriter obtains metadata (e.g., database statistics, query information) to generate a reasoning chain, which includes rewrite proposals and rewritten SQL candidates. This chain is then used to generate the rewritten SQL for further enhancement and validation. Next, ❸ we consolidate the SQL candidates to construct the LLM prompt, and apply global SQL enhancements to find the optimal rewritten SQL. Finally, ❹ we validate the syntactical correctness and semantic equivalence of the rewritten SQL, and ❺ reassess the rationality of the rewritten SQL to decide whether to proceed or rewrite the SQL again based on the rewrite history.
Rewrite Middleware. The toolkit provides three specialized tools to boost LLM rewrite ability. First, we construct a structured knowledge base using a multi-source integration to guide high-quality rewrite proposals (see Section 5.1). Second, a hybrid SQL corrector then validates the syntactical correctness and semantic equivalence of each candidate rewrite, reassessing rationality against rewrite history and triggering further refinement when necessary (see Section 5.2). Third, we design an agent memory buffer that captures key context from previous interactions to reduce communication overhead and LLM hallucination (see Section 5.3).
Query Hint Recommender. The recommender explores the query hints in two steps. ❻ First, we carefully construct a promising query hint base, thereby reducing the hint search space and enabling faster, more accurate hint selection (see Section 6.1). ❼ Second, we leverage LLM agent with data statistics to select appropriate fine-grained hints for the rewritten query to improve its execution performance (see Section 6.2). ❽ Finally, the selected hints are injected into the rewritten SQL to produce the final output.

### v3

# QUITE: A Query Rewrite System Beyond Rules via LLM Agents

@@ 1. Introduction
Inefficient SQL queries remain a long-standing challenge in many database applications, often arising from poorly written queries by inexperienced users or automatically generated systems (cloud; database; extensible). Query rewrite is a promising approach that transforms an original SQL into a semantically equivalent form with better performance (query-rewrite; extensible; wetune). An effective rewritten query can dramatically improve the execution time of a poorly written SQL query by several orders of magnitude (wetune; lr).
Existing query rewrite approaches typically rely on predefined rewrite rules (apache; volcano; pg; lr; lr_demo; querybooster; llm-r2; r-bot). Heuristic-based methods apply these rules in a fixed order derived from practical experience (apache; volcano; pg). However, such fixed orders often fail to generalize across different environments and query patterns. Thus, machine learning (ML)-based methods train models on rewrite histories to predict which rules to use and the order in which to apply them (lr; lr_demo). Recently, some approaches have explored the use of Large Language Models (LLMs) to select rewrite rules and determine their application order based on domain knowledge (llm-r2; r-bot).
@@
However, the above approaches can handle only a small subset of queries and may lead to performance regressions (lr; llm-r2; r-bot). This limitation stems from the intrinsic challenges of rule-based query rewrite. First, the rich features of SQL and the nuances of semantics make it rather difficult to discover and verify new rewrite rules (extensible; hottsql; wetune). Even state-of-the-art rule discovery system (wetune) can only support a narrow subset of operators. Second, fixed rewrite rules rely on pattern matching and therefore are fundamentally unable to optimize unseen or complex query patterns (Leis2015; querybooster; slabcity). Third, many effective rewrite strategies cannot be expressed by rewrite rules (slabcity). For example, the Common Table Expressions (CTE) conversion in Example 1.1 cannot be captured by rewrite rules. This is because such conversion requires explicit control over evaluation order and sub-plan sharing, which is beyond the scope of relational-algebra rewrite rules (graefe1993query; wetune; apache).
Given these limitations, we revisit the query rewrite problem from the perspective of human experts and identify three key capabilities that humans exhibit, but rule-based systems lack. (1) Humans can perform rewrites that follow certain rules but exceed the expressiveness of standard pattern matching, and in some cases, go beyond what can be captured by formal rewrite rules (Example 1.1). (2) Humans can conduct context-aware rewrites by considering the underlying data distributions and cost estimates provided by the query optimizer (Example 1.2). (3) Experts can apply intent-oriented rewrites that capture the underlying goal of the query, which are not easily expressed as explicit rules (Example 1.3).
@@
Query Rewrite Beyond Rewrite Rules. Figure 1 (a) is a simplified query from the DSB benchmark (ding2021dsb). To avoid repeated calculation of nested subqueries, Figure 1 (b) rewrites them using Common Table Expressions (CTEs). This enables the database engine to compute intermediate results once and reuse them, leading to a 37.5 $\times$ reduction in query latency in our experimental setup.
@@
Context-Aware Query Rewrite. In Figure 1 (c), instead of directly joining two large tables (the sale with 10 M rows and the ttime with 1.4 M rows), we first join the sale table with the prod table (100 K rows). This highly selective join would produce an intermediate result of only  100 K rows. Next we join this with ttime table and yield a 2.79 $\times$ speedup over (b) in our experimental setup. Such rewrites, which reduce intermediate size, are rare in rule-based systems and require analyzing data distribution and selectivity.
@@
Query Rewrite based on Query Intent. Figure 2 (a) shows a real-world query from Apache Calcite (apache_calcite_test). Rule-based rewrite cannot optimize this query (Figure 2 (b)). Identifying that the query intent is to sum salaries directly from the EMP table, humans can simplify the original query by eliminating unnecessary joins and sub-queries that do not contribute to the intent of the query, achieving a 2.17 $\times$ execution improvement in our experimental setup.
@@
Although human experts can produce significantly better SQL rewrites than rule-based methods, manual rewriting does not scale well, especially when there are millions of poorly written queries to be optimized in cloud environments. Recently, LLMs have demonstrated nearly human-level semantic and reasoning abilities, making them a natural choice for query rewrite. However, this task remains challenging due to the hallucination problem of LLMs: they may generate responses that are plausible but factually incorrect or semantically invalid. Specifically, there are two fundamental challenges. C1. Ensuring Equivalent Rewrites. Producing equivalent rewrites requires a precise understanding of the rich features of SQL and the nuances of semantics. Even the most powerful LLMs produce rewrites with syntax or semantic errors in nearly 20% of cases (Section 7). C2. Ensuring Optimized Rewrites. Determining whether a rewritten query improves performance depends on multiple factors, including query intent, data distribution, optimizer estimates, system states, and others. These contextual factors are typically unavailable to LLMs, and enabling LLMs to effectively utilize such information remains a non-trivial problem (gptuner; bai-etal-2024-longbench; llmqo).
@@
To address the limitations of rule-based query rewrite, the poor scalability of manual rewrite, and the unreliable nature of LLMs, we propose QUITE, a training-free and feedback-aware system that smartly leverages LLMs to rewrite SQL queries into semantically equivalent forms with significantly improved performance. QUITE supports a broader range of query patterns and rewrite strategies, while maintaining good scalability. It is primarily designed for long-running OLAP workloads, where execution time dominates and the cost of query rewriting is negligible. First, to enable LLMs to capture query data characteristics and leverage database feedback, we design a multi-agent framework controlled by a finite state machine (FSM) to iteratively refine the rewrite process. We decompose the complex rewrite process into a set of subtasks, each handled by a specialized LLM agent to reduce hallucinations and improve rewrite ability (addressing C1, C2). Second, to enhance the ability of LLMs to generate high-quality query rewrites, we provide a rewrite middleware that includes: (1) a structured knowledge base extracted from official documentation and query rewrite discussions on web forums, (2) a hybrid SQL corrector that combines tool-based and LLM-based equivalence checking to ensure query equivalence, Third, we employ a novel hint injection technique to generate better execution plans for rewritten queries. This ensures that query execution exactly follows the optimized plan recommended by LLM agents, avoiding unintended modifications by the query optimizer due to inaccurate cost estimations. Additionally, this technique enables further improvements by incorporating physical-level optimizations (addressing C2).
@@
We extensively evaluate QUITE against state-of-the-art query rewrite systems on widely used benchmarks (e.g., TPC-H, DSB, and Calcite), which shows that our system achieves up to a 21.9% speedup in query execution performance over the best-performing alternative. Moreover, it delivers 24.1% additional rewrites compared to that alternative, covering query cases that previous systems could not handle. In summary, we make the following contributions:
@@
  • We design a multi-agent framework controlled by a finite state machine to enable LLM agents to use external tools and enhance the rewrite process with database feedback (Section 5).
@@ 2. Preliminaries and Related Work
In this section, we first formally define the query rewriting problem and then discuss related work on query hints and LLMs.
@@ 2.1. Query Rewrite
(1) Discovery of new rules. WeTune (wetune) automatically generates and verifies logical plan transformations, but it handles only a limited set of operator types and algebraic rules. Equality saturation techniques (10.1145/3485496) and user-guided systems like QueryBooster (querybooster) still require manual intervention and cover only narrow classes of queries. More recently, GenRewrite (genrewrite) leverages LLMs to describe rewrite rules in natural language (NLR2s), enabling more flexible rule specification. In practice, its effectiveness depends on the coverage of available equivalence checks, and validating a large number of candidate rewrites through execution can introduce substantial overhead (genrewrite; slabcity).
(2) Effective use of existing rules. LearnedRewrite (lr; lr_demo) uses Monte Carlo Tree Search (MCTS) with learned cost models to explore the space of rule applications, but its performance depends on the accuracy of cost model estimation. LLM-R2 (llm-r2) retrieves similar past rewrites using a pre-trained selector and employs them as in-context examples for rule selection. R-Bot (r-bot) adopts retrieval-augmented generation to recommend rewrite rules through evidence retrieval and step-by-step reasoning. These approaches largely retrieve and apply rewrite knowledge based on query structure and then select from predefined rule sets. Consequently, they may not fully capture dataset-dependent factors such as database scale or column statistics, and their rule-centric formulation can miss effective expert strategies that are hard to express as explicit rewrite rules.
@@ 2.2. Query Hints
Query hints are SQL extensions that provide instructions to the database’s query engine to influence the selection of execution plans (bao; autosteer; fastgres; steeringqo). These hints allow users to exert precise control over how the database optimizer chooses physical operators and access methods. As shown below for pg_hint_plan extension (pg_hint_plan), the hint /*+ HashJoin(employees departments) */ forces the query engine to use a hash join for the $employees\bowtie departments$ operation. This hint can override the optimizer’s default cost-based plan selection.
/*+ HashJoin(employees departments) */
@@
Prior works on query hint selection (bao; autosteer; fastgres) apply boolean hints to entire queries, which often leads to suboptimal performance because different operators within a query may require different hints. Proto-X (zhang2024holon) improves upon this by assigning hints at the operator level. However, it operates in an offline setting that requires exploring multiple configurations and repeatedly executing the complete workload, whereas query rewriting does not require repeated workload execution. UniTune (zhang2023unified) supports query rewriting through LearnedRewrite (lr) but does not incorporate hints. In contrast, QUITE employs pg_hint_plan to enable fine-grained, operator-level control (e.g., selection of join algorithms and adjustment of cardinalities) and jointly optimizes both query rewriting and hint selection.
@@ 2.3. LLM and LLM Agents
LLM for Database Tasks. LLMs are increasingly applied to database tasks text-to-SQL (text2sql2; codes; omnisql; xiyansql), SQL workload generation (SQLBarber; lao2025sqlbarberleveraginglargelanguage), and database optimization (gptuner; llmidx; gptuner-demo; llmqo; gptuner-record), with recent work focusing on query optimization (llm-r2; r-bot; genrewrite). Although these methods show that LLMs can understand and modify SQL, the inherent complexity of query rewrite presents substantial challenges.
LLM Agents. The primary advantage of LLM agents over LLMs is their ability to collect feedback from the execution environment (e.g., database execution metrics) and adapt their subsequent reasoning and outputs accordingly (wu2025agenticreasoningreasoningllms). They have proven particularly effective in code generation (liu2023your; yang2024sweagentagentcomputerinterfacesenable), automated diagnostics (d-bot; mdagent; thompson2023large), and knowledge management (d-bot; asai2024self). Moreover, collaborative deployments of multiple LLM agents often outperform single‑agent systems on complex reasoning and planning tasks (multi1; multi2; multi3; multi4). QUITE applies this multi-agent methodology to query rewrite by gathering DBMS feedback and adjusting each query candidate accordingly.
@@ 3. System Overview
We present the architecture of QUITE  as shown in Figure 3. It comprises three components: LLM Agent-Based Query Rewriter, Rewrite Middleware, and Query Hint Recommender.
LLM Agent-Based Query Rewriter. The rewriter uses a training-free and feedback-aware LLM agent-based FSM to rewrite queries in multiple stages. First, ❶ the user provides the input SQL along with the corresponding database configuration (e.g., DBMS type, host address). Second, ❷ the rewriter obtains metadata (e.g., database statistics, query information) to generate a reasoning chain, which includes rewrite proposals and rewritten SQL candidates. This chain is then used to generate the rewritten SQL for further enhancement and validation. Next, ❸ we consolidate the SQL candidates to construct the LLM prompt, and apply global SQL enhancements to find the optimal rewritten SQL. Finally, ❹ we validate the syntax correctness and semantic equivalence of the rewritten SQL, and ❺ verify the correctness and effectiveness of the rewritten query to decide whether to proceed to Step ❻ or further rewrite this query based on the rewrite history.
Rewrite Middleware. The toolkit provides three specialized tools to boost LLM rewrite ability. First, we construct a structured knowledge base using a multi-source integration to guide high-quality rewrite proposals (Section 5.1). Second, a hybrid SQL corrector then validates the syntactical correctness and semantic equivalence of each candidate rewrite, using rewrite history to trigger further refinement when necessary. (Section 5.2). Third, we design an agent memory buffer that captures key context from previous interactions to reduce communication overhead and LLM hallucination (Section 5.3).
Query Hint Recommender. The recommender explores the query hints in two steps. ❻ First, we construct a promising query hint base, thereby reducing the hint search space and enabling faster, more accurate hint selection (Section 6.1). ❼ Second, we leverage an LLM agent with data statistics to select appropriate fine-grained hints for the rewritten query to improve its execution performance (Section 6.2). ❽ Finally, the selected hints are injected into the rewritten SQL to produce the final output.
Figure 4. Reasoning Chain of MDP-based Reasoning Agent